OpenStudy (anonymous):

Find the range of the function f(x) = 2x + 3 for the domain {0,1,2}.

jimthompson5910 (jim_thompson5910):

Plug in x = 0 to get f(x) = 2x+3 f(0) = 2(0)+3 f(0) = 3 So when x = 0, f(x) = 3 This means that 3 is part of the range To find the other numbers in the range, repeat what is shown above but do it for x = 1 and x = 2